# Unveiling Challenges for LLMs in Enterprise Data Engineering

### v1

@@ 6. The Knowledge Challenge
Since LLMs are primarily pre-trained on public data, they lack information about a company’s internal business processes and policies, as well as about its proprietary tools and systems (Kayali et al., 2025). In this section, we analyze how this lack of company-specific knowledge in LLMs affects their accuracy on enterprise data engineering tasks.
@@ 6.1. Knowledge Challenges
We identify two main categories of challenges: (1) the lack of knowledge that is not available to LLMs but does exist in documentation or other company-internal sources, and (2) company-specific extensions of databases that are typically not documented at all.
C9: Proprietary but available knowledge. Enterprises often use proprietary tools and systems to address data engineering challenges. In contrast to well-established technologies like SQL, there is typically much less documentation about them available on the web. Therefore, it is reasonable to assume that LLMs trained on public data have little parametric knowledge about these tools and systems. One example is the domain-specific query language SIGNAL (Kampik et al., 2023), which SAP provides to its customers for exploring data about business processes. SIGNAL resembles SQL in many aspects but also includes domain-specific features and syntax differences tailored towards process mining. While the language itself is well-specified on its public documentation page, it has a much smaller user base than SQL and, therefore, a smaller online footprint regarding help pages, Q&A threads, and blog posts, resulting in limited understanding. Nevertheless, SIGNAL is frequently used by data engineers and must thus be well-supported by LLMs, for example, to translate natural language requests into SIGNAL queries.
C10: Proprietary and unavailable knowledge. Enterprise systems like those from SAP support company-specific changes and extensions to customize the system for individual customers. For example, customers can use hooks to add custom business logic. On the data level, customization means extending the database schema by adding customer-specific columns to existing tables or even additional tables to the customer namespace. In our analysis of real-world systems from SAP, we have come across thousands of such customer-defined tables. Since these changes are highly company-specific, they are typically not documented publicly—if they are documented at all. Moreover, customers sometimes “misuse” existing attributes of the standard SAP schema for different purposes. Such digressions from the public documentation pose significant challenges for data engineering with LLMs as well.
@@ 6.2. Experiments & Results
To analyze how the challenges C9-C10 affect data engineering with LLMs, we perform two experiments. (1) First, we apply LLMs to the task of translating natural language requests into the proprietary query language SIGNAL. As mentioned in C9, the specifics of SIGNAL provide a good example for proprietary knowledge that is available primarily in company-internal documentation. (2) In a second experiment, we exercise the case where no knowledge is available to enrich the context of LLMs, not even in internal documentation (C10). In this experiment, we want to determine to what extent LLMs understand customer-specific extensions to the standard SAP schema. With a simple probing task, we demonstrate how LLMs struggle with customizations and highlight their strong bias towards publicly documented parts of the schema.
Exp. 9: Text-to-SIGNAL. In our first experiment, we use LLMs to translate natural language requests into the proprietary query language SIGNAL, as shown in Figure 2 (right).
@@
For this experiment, we use a set of 200 randomly-sampled pairs of natural language requests and SIGNAL queries from a larger dataset in use at SAP. Given a natural language request, like Retrieve the number of unique invoices from ’defaultview-255’, we prompt the LLMs to generate a corresponding SIGNAL query. We compare three different approaches: First, we prompt the LLMs only with a short instruction explaining the task (Zero-shot). Next, we add three fixed representative example queries that we hand-picked from the SIGNAL documentation (+ Examples). Finally, we include both the examples as well as the full documentation of the SIGNAL language in the prompt (+ Documentation). For all three scenarios, we manually tune our prompts to alleviate obvious mistakes and hint at the differences between SIGNAL and SQL. To evaluate the generated queries, we compute the execution accuracy by executing the ground truth and predicted query and comparing their results.
@@ Results.
Figure 11 shows the results of our Text-to-SIGNAL experiment (left bar groups). We can clearly see that only a small fraction (up to ) of the generated SIGNAL queries is correct. Moreover, we find that adding example queries to the prompt brings the biggest improvements in accuracy. By contrast, additionally including the full documentation of the SIGNAL language in the prompt improves the accuracy only for some models, indicating that simply providing the required documentation during inference is not a viable solution. Finally, a comparison with the results on popular Text-to-SQL benchmarks (Yu et al., 2018; Li et al., 2023) (Figure 11 right) shows that the accuracy for Text-to-SIGNAL is substantially lower.
@@
Why are examples and documentation not helping? To better understand why exactly the models fail, we conduct an error analysis that categorizes the queries into those that execute with (1) correct or (2) incorrect results, and those that fail to execute because of (3) syntactic or (4) semantic errors. We focus on GPT-4o, the only model that could correctly translate a non-negligible number of requests in all three configurations. Figure 12 (left) shows that most of the generated queries fail to execute because of syntactic or semantic errors, with the largest number of failures caused by syntax errors. Adding example queries and documentation to the prompt causes the number of syntax errors to decrease. Nevertheless, we also see that the number of queries that contain semantic errors or return incorrect results increases as we add examples and documentation. A closer inspection of the generated queries reveals that the models often confuse parts of the documentation, like the column names used in explanations, with the request they have to translate. As a result, the number of fully correct queries is far from satisfactory even when including representative examples and the full SIGNAL documentation in the prompt.
@@
A bias towards SQL. Another interesting effect we observe is that the models tend to generate SQL-specific syntax and ignore the specifics of SIGNAL. To further dive into the root causes, we manually analyze the syntax errors generated by GPT-4o in the (+ Documentation) setting and group them into three categories: (1) incorrect use of GROUP BY and ORDER BY, which in SIGNAL requires numerical indices instead of column names (unlike in SQL), (2) use of invalid characters, like an asterisk (*) in a count statement which exists in SQL but not in SIGNAL, and (3) incorrect structure of the overall statement. As shown in Figure 12 (right), we find that even when explicitly including rules in the prompt to avoid such SQL-related syntax errors, the models still generate erroneous SIGNAL queries. This behavior indicates a strong bias of LLMs towards SQL, which may be caused by its much higher prominence in their pre-training data. Similar problems could also exist in other scenarios where enterprise data or tasks differ only slightly from public knowledge, like currency conversion or complex data types.
Exp. 10: Schema customizations. In our next experiment, we analyze how LLMs perform when a task requires company-specific knowledge for which there is little to no explicit documentation.
Setup.
For this experiment, we use a scenario with customizations of the standard SAP schema and analyze to which extent the LLMs’ internal parametric knowledge already captures this knowledge. To see to which extent the parametric knowledge of LLMs is able to understand customized tables without introducing further task-specific difficulties, we choose a simple task setup: Given an SAP table name, we instruct the model to generate a list of all columns that belong to that table. That way, we can see to which extent LLMs know the table columns. In our analysis, we differentiate between columns that are part of the standard SAP schema and those that were defined by the customer. Our corpus consists of 2,000 tables from a real-world SAP system, of which 1,000 are part of the standard SAP schema and 1,000 were defined by the customer. We evaluate the predictions by computing F1 scores that measure the overlap between the predicted and true columns of each table.
Results.
Figure 13 shows the results broken down by the frequency of how often columns appear in the schema. As expected, the parametric knowledge about customer-defined columns is severely limited to certain columns like MANDT that must appear in most customer-defined tables. In contrast, the LLMs achieve higher F1 scores for columns from standard SAP tables. Still, we see that their parametric knowledge is strongly biased towards columns that appear in many tables and are thus more prominent in public documentation. Importantly, our introductory experiment (Figure 1 right) shows that this lack of enterprise-specific knowledge about customized columns severely impacts performance on downstream tasks like column type annotation.
@@ 6.3. Discussion
Our experiments show that the lack of enterprise knowledge affects the performance of LLMs. Even when including hand-picked examples or the full documentation in the prompt, LLMs are unable to actualize this information to solve enterprise tasks like Text-to-SIGNAL. An interesting finding here is that biases in the LLMs’ knowledge towards public sources, like their deep understanding of SQL, are hard to overcome even with explicit instructions, making LLMs difficult to adapt to a wide variety of specialized enterprise use cases without more heavy-weight solutions like fine-tuning that cause additional overheads.
@@
As discussed before, our experiments show that simple out-of-the-box approaches like few-shot prompting and including relevant documentation in the prompt are insufficient. While we believe that more involved approaches, like fine-tuning models for specific enterprise use cases, could help them acquire the necessary domain knowledge, they require extensive work, for example to label the necessary data for fine-tuning. As such, we believe that future research should explore the direction of self-supervised pre-training of foundation models on large enterprise corpora, specifically on enterprise data and documentation. This could allow them to acquire a broad base of enterprise knowledge that enables them to solve complex enterprise tasks out of the box, which is an idea that has very recently gained some attraction in industry.444https://www.databricks.com/company/newsroom/press-releases/databricks-and-anthropic-sign-landmark-deal-bring-claude-models
@@ 7. Costs At Enterprise Scale
As described in Section 4, enterprise databases are substantially larger than typical data engineering benchmark datasets, with SAP databases often reaching multiple terabytes in size. Table 6 shows how this large size translates into high LLM costs. Processing one gigabyte of tabular data from our SAPCTA corpus already costs about USD 71. Note that this cost does not include generating any output tokens, which are substantially more expensive than input tokens, as shown in Table 1. Nevertheless, looking only at the input token costs already provides a rough estimate of the expected high costs. For example, processing the entire customer database with GPT-4o-Mini (OpenAI’s cheapest model) would cost USD 463T, and using OpenAI’s o1 model would cost USD 46M, as shown in Table 6. These high costs clearly prohibit using LLMs on large fractions of the enterprise data. This is especially problematic as our experiments in Sections 4-6 have shown that cheaper models like GPT-4o-Mini often underperform more expensive models.
@@
As a second cost driver, we want to highlight the algorithmic complexity of the ways in which many data engineering tasks are currently approached with LLMs. One example is the task of entity matching, which current approaches address by using LLMs to compare pairs of entities (Peeters and Bizer, 2023; Li et al., 2024). This approach has an algorithmic complexity of , where and are the numbers of entities in each table, making it intractable even for medium-sized tables. For example, matching just 1,000 payments to 1,000 invoices already costs USD 1,462, and matching 10,000 payments to 10,000 invoices would cost USD 146,165. Future work should therefore put a strong emphasis on reducing the costs of LLM-based approaches, for example by exploring task formulations with lower complexities (Wang et al., 2025), using LLMs together with smaller fine-tuned models (Liu et al., 2024b; Zhang et al., 2024b), or combining them with conventional approaches such as blocking (Zhang et al., 2020b).
@@ 8. Conclusion
In this paper, we systematically analyzed the challenges involved in applying LLMs to real-world enterprise data engineering scenarios. Through experiments on a diverse set of tasks, we have shown how their effectiveness is constrained by enterprise-specific complexities along the dimensions of data, tasks, and knowledge, significantly impacting the reliability of LLM-driven automation in enterprises. We hope that our insights and learnings provide a helpful guide and inspiration for future efforts to make LLMs viable for enterprise data engineering and support their adoption in industry.
@@ Acknowledgements.
This work has been supported by the BMBF and the state of Hesse as part of the NHR Program and the HMWK cluster project 3AI. It was also partially funded by the LOEWE Spitzenprofessur of the state of Hesse. We also thank DFKI Darmstadt, SAP, and hessian.AI as well as Dr. Alexey Streltsov, Matthias Urban, and Furkan Karakocaoglu for their support.References
@@
- Bodensohn et al. (2024a) Jan-Micha Bodensohn, Ulf Brackmann, Liane Vogel, Anupam Sanghi, and Carsten Binnig. 2024a. Automating Enterprise Data Engineering with LLMs. In NeurIPS 2024 Third Table Representation Learning Workshop. https://openreview.net/forum?id=m85fYEJtDc
@@
- He et al. (2025) Xinyi He, Yihao Liu, Mengyu Zhou, et al. 2025. TableLoRA: Low-rank Adaptation on Table Structure Understanding for Large Language Models. https://doi.org/10.48550/arXiv.2503.04396 arXiv:2503.04396 [cs].
@@
- Liu et al. (2025) Chunwei Liu, Matthew Russo, Michael Cafarella, et al. 2025. Palimpzest: Optimizing AI-Powered Analytics with Declarative Query Processing. In 15th Annual Conference on Innovative Data Systems Research.
- Liu et al. (2024a) Nelson F. Liu, Kevin Lin, John Hewitt, et al. 2024a. Lost in the Middle: How Language Models Use Long Contexts. Trans. Assoc. Comput. Linguistics 12 (2024), 157–173. https://doi.org/10.1162/TACL_A_00638
@@
- Liu et al. (2024b) Yurong Liu, Eduardo Pena, Aecio Santos, Eden Wu, and Juliana Freire. 2024b. Magneto: Combining Small and Large Language Models for Schema Matching. https://doi.org/10.48550/arXiv.2412.08194 arXiv:2412.08194 [cs].
@@
- OpenAI (2024b) OpenAI. 2024b. Learning to Reason with LLMs. https://openai.com/index/learning-to-reason-with-llms/.
@@
- Zhang et al. (2020a) Dan Zhang, Madelon Hulsebos, Yoshihiko Suhara, Çağatay Demiralp, Jinfeng Li, and Wang-Chiew Tan. 2020a. Sato: contextual semantic type detection in tables. Proceedings of the VLDB Endowment 13, 12 (Aug. 2020), 1835–1848. https://doi.org/10.14778/3407790.3407793
@@
- Zhang et al. (2020b) Wei Zhang, Hao Wei, Bunyamin Sisman, Xin Luna Dong, Christos Faloutsos, and Davd Page. 2020b. AutoBlock: A Hands-off Blocking Framework for Entity Matching. In Proceedings of the 13th International Conference on Web Search and Data Mining (WSDM ’20). Association for Computing Machinery, New York, NY, USA, 744–752. https://doi.org/10.1145/3336191.3371813

### v2

@@ 6. The Knowledge Challenge
Since LLMs are primarily trained on public data, they lack information about a company’s internal business processes and policies, as well as about its proprietary tools and systems (Kayali et al., 2025). In this section, we analyze how the lack of company-specific knowledge in LLMs affects their accuracy on data engineering tasks.
@@ 6.1. Knowledge Challenges
We identify two main categories of challenges: the lack of knowledge that is not available to LLMs but does exist in documentation and other company-internal sources, and company-specific extensions of databases that are typically not documented at all.
C9: Proprietary but available knowledge. Enterprises often use proprietary tools and systems to solve data engineering challenges. In contrast to well-established technologies like SQL, there is typically much less documentation about them available on the web. Therefore, it is reasonable to assume that LLMs trained on public data have little parametric knowledge about these tools and systems. One example is the domain-specific query language SIGNAL (Kampik et al., 2023), which SAP provides to its customers for exploring data about business processes. SIGNAL resembles SQL in many aspects but also includes domain-specific features and syntax differences tailored towards process mining. While the language itself is well-specified on its public documentation page, it has a much smaller user base than SQL and, therefore, a smaller online footprint in terms of help pages, Q&A threads, and blog posts. Nevertheless, SIGNAL is frequently used by data engineers and must thus be well-supported by LLMs, for example, to translate natural language requests into SIGNAL queries.
C10: Proprietary and unavailable knowledge. Enterprise systems like those from SAP support company-specific changes and extensions to customize the system for individual customers. For example, customers can use hooks to add custom business logic. On the data level, customization means extending the database schema by adding customer-specific columns to existing tables or even additional tables to the customer namespace. In our analysis of real-world systems at SAP, we have come across thousands of such customer-defined tables. Since these changes are highly company-specific, they are typically not documented publicly—if they are documented at all. Moreover, customers sometimes “misuse” existing attributes of the standard SAP schema for different purposes. Such digressions from the public documentation pose significant challenges for data engineering with LLMs as well.
@@ 6.2. Experiments & Results
To analyze how challenge C9 affects data engineering with LLMs, we use LLMs to translate natural language requests into the domain-specific query language SIGNAL. The specifics of SIGNAL provide a good example for the proprietary knowledge primarily available in company-internal documentation. For challenge C10, we refer to Figure 1 as well as an additional experiment in Appendix A.5 of our extended technical report.
Exp. 10: Text-to-SIGNAL. In this experiment, we use LLMs to translate natural language requests into the proprietary query language SIGNAL, as shown in Figure 2 (right).
@@
We experiment on a set of 200 randomly-sampled pairs of natural language requests and SIGNAL queries from a larger dataset in use at SAP. Given a request like Retrieve the number of unique invoices from ’defaultview-255’, we prompt the LLM to generate a corresponding SIGNAL query. We compare three approaches: First, we provide the LLM only with a short instruction explaining the task (Zero-shot). Next, we add three fixed example queries that were hand-picked from the SIGNAL documentation (+ Examples). Finally, we include both the examples as well as the documentation of the SIGNAL language in the prompt (+ Documentation). We manually tune our prompts in all scenarios to alleviate obvious mistakes and hint at the differences between SIGNAL and SQL. To evaluate the generated queries, we compute the execution accuracy by executing the ground truth and predicted query and comparing their results.
@@ Results.
Figure 10 (left) shows the results of this experiment. Only a small fraction of the generated SIGNAL queries are correct. Adding example queries to the prompt brings the biggest improvements in accuracy. By contrast, additionally including the documentation of the SIGNAL language in the prompt improves the accuracy only for some models, indicating that simply providing the required documentation during inference is not a viable solution. Finally, a comparison with the results on popular Text-to-SQL benchmarks (Yu et al., 2018; Li et al., 2023) (Figure 10 right) shows that the out-of-the-box accuracy for Text-to-SIGNAL is substantially lower.
Why are examples and documentation not helping?
@@
To better understand why exactly the models fail, we conduct an error analysis that categorizes the queries into those that execute with correct or incorrect results, and those that fail to execute because of syntactic or semantic errors. We focus on GPT-4o, the only model that correctly translated a non-negligible number of requests in all three configurations. Figure 11 (left) shows that most of the generated queries fail to execute because of syntactic or semantic errors, with the largest number of failures caused by syntax errors. Adding example queries and documentation to the prompt causes the number of syntax errors to decrease. However, the number of queries that contain semantic errors or return incorrect results increases as we add examples and documentation. A closer inspection of the generated queries reveals that the models often confuse parts of the documentation, like the column names used in explanations, with the request they have to translate. As a result, the number of fully correct queries is far from satisfactory even when including hand-picked examples and the documentation in the prompt.
A bias towards SQL.
@@
Another interesting effect is that the models tend to generate SQL-specific syntax and ignore the specifics of SIGNAL. To further dive into the root causes, we manually analyze the 75 syntax errors generated by GPT-4o in the (+ Documentation) setting and group them into three categories: the incorrect use of GROUP BY and ORDER BY, which in SIGNAL requires numerical indices instead of column names, the use of invalid characters, like an asterisk (*) in a count statement which exists in SQL but not in SIGNAL, and the incorrect structure of the overall statement. As shown in Figure 11 (right), we find that even when explicitly including rules in the prompt to avoid such SQL-related syntax errors, the models still generate erroneous SIGNAL queries. This behavior indicates a strong bias of LLMs towards SQL, which may be caused by its much higher prominence in their training data. Similar problems could also exist in other scenarios where enterprise data and tasks differ only slightly from public knowledge, like currency conversion or company-specific data types.
@@ 6.3. Discussion
Our experiments show that the lack of enterprise-specific knowledge affects the performance of LLMs. Even when including hand-picked examples and the full documentation in the prompt, LLMs are unable to actualize this information to solve enterprise tasks like Text-to-SIGNAL. An interesting finding here is that biases in the models’ knowledge towards public sources, like their understanding of SQL, are hard to overcome even with explicit instructions, making LLMs difficult to adapt without more heavy-weight solutions like fine-tuning that cause additional overheads.
@@
Our results demonstrate that simple out-of-the-box approaches like few-shot prompting and including relevant documentation in the prompt are insufficient. While we believe that more involved approaches, like fine-tuning LLMs for specific enterprise use cases, could help them acquire the necessary domain knowledge, these approaches require extensive work, for example to label the necessary data for fine-tuning for each individual task. Moreover, even self-supervised (task-agnostic) pre-training on enterprise corpora comes with important challenges: enterprise data often contains valuable and sensitive information, which raises significant privacy and copyright concerns. Therefore, we believe that future research should explore controlled alternatives, such as working with synthetic or anonymized corpora that reflect enterprise data characteristics (Sanghi and Haritsa, 2023). Despite all these challenges, we believe that LLMs tuned to enterprise settings are a promising direction and ideas in that direction have recently gained traction in industry to adapt foundation models for enterprise environments.555https://www.databricks.com/company/newsroom/press-releases/databricks-and-anthropic-sign-landmark-deal-bring-claude-models (last accessed November 4, 2025)
@@ 7. Costs At Enterprise Scale
As described in Section 4, enterprise databases are substantially larger than typical data engineering benchmark datasets, with SAP databases often reaching multiple terabytes in size. Table 7 shows how this large size translates into high LLM costs. Processing one gigabyte of tabular data from our SAPCTA corpus with GPT-4o-Mini (OpenAI’s cheapest model) already costs about USD 71. Note that this cost does not include generating any output tokens, which are substantially more expensive than input tokens, as shown in Table 1. Nevertheless, the input token costs already provide a rough estimate of the expected high costs. For example, processing the entire customer database with GPT-4o-Mini would cost USD 463T, and using the o1 reasoning model would cost USD 46M. These high costs clearly prohibit using LLMs on large fractions of the enterprise data. This is especially problematic as our experiments in Sections 4-6 have shown that cheaper models like GPT-4o-Mini often underperform more expensive models.
@@
As a second cost driver, we want to highlight the algorithmic complexity of the ways in which many data engineering tasks are currently approached with LLMs. One example is the task of entity matching, which current approaches address by using LLMs to compare pairs of entities (Peeters and Bizer, 2023; Li et al., 2024). This approach has an algorithmic complexity of , where and are the numbers of entities in each table, making it intractable even for medium-sized tables. For example, matching just 1,000 payments to 1,000 invoices already costs USD 1,462, and matching 10,000 payments to 10,000 invoices would cost USD 146,165. Future work should therefore focus on reducing the cost of LLM-based approaches by rethinking task formulations and combining LLMs with smaller fine-tuned models (Liu et al., 2025a; Zhang et al., 2024b) and other techniques. For example, for entity matching, Wang et al. (2025) propose a task formulation where, given a row from one table, the LLM is asked to identify all matching rows from the other table in one step. This reduces the complexity to if the second table fits in context, and to when batching is used with batch size , offering a substantial reduction in cost without compromising accuracy.
@@ 8. Conclusion
In this paper, we systematically analyzed the challenges involved in applying LLMs to real-world data engineering scenarios in enterprises. Through experiments with multiple LLMs on a diverse set of tasks, we have shown that the effectiveness of these models is substantially affected by enterprise-specific complexities along the dimensions of data, tasks, and knowledge. While the absolute numbers of the observed effects may change with more experiments and newer models, the large gap between enterprise and public scenarios shows the clear need for future research. On the same note, we see a strong need for enterprise-specific LLMs, with many recent activities emerging from research and industry. We hope that our insights and learnings provide a helpful guide and inspiration for future efforts to make LLMs viable for enterprise data engineering and support their adoption in industry.
@@ Acknowledgements.
This work has been supported by the BMBF and the state of Hesse as part of the NHR program and the HMWK cluster project 3AI. It was also partially funded by the LOEWE Spitzenprofessur of the state of Hesse and has benefited from early stages of funding from the Deutsche Forschungsgemeinschaft (DFG) under Germany’s Excellence Strategy (EXC-3057); funding will begin in 2026. We thank DFKI Darmstadt, SAP, and hessian.AI as well as Alexey Streltsov, Matthias Urban, and Furkan Karakocaoglu for their support.References
@@
- Bodensohn et al. (2024a) Jan-Micha Bodensohn, Ulf Brackmann, Liane Vogel, Anupam Sanghi, and Carsten Binnig. 2024a. Automating Enterprise Data Engineering with LLMs. In NeurIPS 2024 Third Table Representation Learning Workshop. https://openreview.net/pdf?id=m85fYEJtDc
@@
- He et al. (2025) Xinyi He, Yihao Liu, Mengyu Zhou, et al. 2025. TableLoRA: Low-rank Adaptation on Table Structure Understanding for Large Language Models. In Proceedings of the 63rd Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers), ACL 2025, Vienna, Austria, July 27 - August 1, 2025, Wanxiang Che, Joyce Nabende, Ekaterina Shutova, and Mohammad Taher Pilehvar (Eds.). Association for Computational Linguistics, 22376–22391. https://aclanthology.org/2025.acl-long.1090/
@@
- Liu et al. (2025b) Chunwei Liu, Matthew Russo, Michael Cafarella, et al. 2025b. Palimpzest: Optimizing AI-Powered Analytics with Declarative Query Processing. In 15th Annual Conference on Innovative Data Systems Research.
- Liu et al. (2024) Nelson F. Liu, Kevin Lin, John Hewitt, et al. 2024. Lost in the Middle: How Language Models Use Long Contexts. Trans. Assoc. Comput. Linguistics 12 (2024), 157–173. https://doi.org/10.1162/TACL_A_00638
@@
- Liu et al. (2025a) Yurong Liu, Eduardo Peña, Aécio S. R. Santos, Eden Wu, and Juliana Freire. 2025a. Magneto: Combining Small and Large Language Models for Schema Matching. Proc. VLDB Endow. 18, 8 (2025), 2681–2694. https://www.vldb.org/pvldb/vol18/p2681-freire.pdf
@@
- OpenAI (2024b) OpenAI. 2024b. Learning to Reason with LLMs. https://openai.com/index/learning-to-reason-with-llms/. Accessed 18-02-2025.
@@
- Papicchio et al. (2025) Simone Papicchio, Paolo Papotti, and Luca Cagliero. 2025. QATCH: Automatic Evaluation of SQL-Centric Tasks on Proprietary Data. ACM Trans. Intell. Syst. Technol. 16, 2 (2025), 45:1–45:26. https://doi.org/10.1145/3712704
@@
- Sanghi and Haritsa (2023) Anupam Sanghi and Jayant R. Haritsa. 2023. Synthetic Data Generation for Enterprise DBMS. In 2023 IEEE 39th International Conference on Data Engineering (ICDE). 3585–3588. https://doi.org/10.1109/ICDE55515.2023.00274
@@
- Zhang et al. (2020) Dan Zhang, Madelon Hulsebos, Yoshihiko Suhara, Çağatay Demiralp, Jinfeng Li, and Wang-Chiew Tan. 2020. Sato: contextual semantic type detection in tables. Proceedings of the VLDB Endowment 13, 12 (Aug. 2020), 1835–1848. https://doi.org/10.14778/3407790.3407793